# How Linguistics Learned to Stop Worrying and Love the Language Models

### v1

@@ 4.6 Upshots for linguistics beyond language structure
In much of this piece, we have focused on the upshot of LMs for linguistic structure, particularly syntax. We did so for two main and related reasons: first, because of the centrality of structure for Big Questions about language over the last 60+ years; and, second, because mastery of linguistic form emerged in models earlier than other high-level abilities (Mahowald et al.,, 2024), seemingly largely from the pretraining objective. Taken together, these points have led to important insights about human language learning and processing.
But LMs have made striking gains in domains like reasoning, logic, and long-form dialog, even since the introduction of ChatGPT in 2022. These abilities seem to emerge not just from pretraining, but from techniques like supervised fine-tuning, instruction-tuning, and RLHF whereby models are given specific feedback so as to make them more aligned with human behavior on specific tasks (Bai et al.,, 2022; Achiam et al.,, 2023). The importance of inference-time compute—whereby models learn in-context and perform additional computation during the process of generating text—has also become apparent for tasks like reasoning (OpenAI et al.,, 2024). It is becoming increasingly difficult to write off these successes as mere parroting.
As such, just as it is fruitful to explore what inductive biases and representational mechanisms underlie the emergence of syntax in LMs, it is increasingly fruitful to study what inductive biases and structures are necessary and/or sufficient for other abilities in LMs like reasoning, theory of mind, planning, and other aspects of higher-level cognition. Many of our same arguments hold in these domains as well: the upshot of the modern view of learning, the contravariance principle, the role of the statistical tradition, the value of real pattern accounts.
Moreover, as LMs continue to develop, they increasingly part of our speech communities. Millions of humans regularly interact with chatbots, in some cases carrying on long extended dialogs with chatbots playing various roles. And, increasingly, text read by humans is written or co-written by AI. How will the presence of AI change language? How should we think about dialogs between human and non-human interlocutors? What social consequences will emerge from this new paradigm? Linguistics as a field, particularly sociocultural linguistics and related disciplines, is well-positioned to be at the forefront of answering these questions by drawing on an already large existing body of work and theory.
@@ 5 Conclusion
Rather than resisting this development, we contend it offers a golden opportunity for linguistics by enabling new kinds of research and opening up vistas of new hypotheses, hypotheses that we think both Norm and Claudette will be excited about. Some of these questions will be new, like what kinds of artificial neural architectures best capture language, what kinds of biases they hold, and what the sources of those biases are. But we also have highlighted ways in which these methods can make progress on some of the oldest and most venerable questions in linguistics: what must be true of the input data for certain structures to be learned in principle, what are the constraints on what languages are possible, how does the form of language relate to its function.
Our response to all this is not to throw out linguistic structure and theory. In response to Piantadosi,’s claim that language models refute Chomsky, Kodner et al., (2023) wrote a piece skeptical of the role of LMs in linguistic theory called “Why linguistics will thrive in the 21st century”. We would go a step further and argue that, with the 21st century a quarter over, linguistics is thriving. But we think it’s doing so not in spite of language models, but in part because of them and the opportunities they open up for new research directions and the light they shed on old questions.
Therefore, we reject the dichotomy in the discourse around Piantadosi,’s (2023) paper, which often seems to pit language models against linguistics. Linguistic structure is real and important even if language models learn those structures emergently in a complex statistical way. Conversely, we think language models can point the way to new ways of thinking about the learning, processing, and fundamental nature of language.
@@
  • Much of the field of linguistics has assumed that the form of language must be explained in terms of a discrete grammar formalism that constrains the forms of possible languages, and that this grammar formalism must represent innate human constraints on language. The success of machine learning models should orient us away from this paradigm, because it shows that this approach is not necessary to characterize learning, and it opens up a new universe of statistical, quantitative, and functional theories to constrain the forms of possible languages and explain why only humans have it, while at the same time providing the tools to test those theories.
@@
  • Relatedly, the idea that the formal structure of linguistic competence should be the main focus of theoretical linguistics, which precedes other forms of study (Chomsky,, 1965), is no longer tenable—language models reveal and allow us to characterize new dimensions of language such as how corpora of text express world knowledge, how the structure of usage supports learning or doesn’t, and how fundamental information-processing constraints shape the way that language is represented in brains and machines.
@@
  • The success of machine learning models shows us how language can be represented in ways that are graded, probabilistic, and fuzzy, which should move us away from an insistence on discrete categorical frameworks. They also reveal potential soft constraints on the structure of language, as reviewed in Section 4.4.
@@
  • The ontological basis for linguistic structure need not lie in an innate genetic endowment: linguistic structure is a real pattern, just as real and worthy of study whether it lies innate in the human genome or whether it is learned entirely through inductive statistical learning with domain-general biases. Linguistic phenomena do not become less interesting when they are learnable in this way, they become more interesting.
@@
More broadly, language models show that linguistics needs the rest of science. We will not answer the most fundamental questions about language by focusing only on language, using only methods that are tailor-made for linguistic analysis. Just as neural networks revealed principles that are now fundamental in our understanding of vision, they are now in a position to revolutionize our understanding of language. Just as physics was revolutionized by thermodynamic ideas which arose from the ad-hoc analysis of complex machines, linguistics stands to benefit from new ways of thinking about computation arising from the analysis of neural networks. There is now massive intellectual activity spurred by language models in cognitive science, philosophy, physics, statistical learning theory, and information theory, and the science of language can draw ideas, methods, and inspiration from all of this by maintaining a spirit of deep, curious, open-minded engagement and integration. Human language is in many ways unique as a natural phenomenon, and this means that the science of language should integrate with other fields of science more, in order to find deeper unities and to delineate and explain how and why language is special.
@@
In particular, we are optimistic about revising what Stabler, (1983) said in doubting that “any close connections between linguistic theory and biology will be forthcoming”. He continues: “it is a crucial advantage of the computational approach that it has a functionalist vocabulary that does not require a type reduction to physicalistic concepts […] since this seems to be required in linguistics and in psychological accounts of language processing and visual perception”. We agree that there is a crucial advantage to the functionalist vocabulary used in linguistic theory, but by analogy to recent work in vision, we are optimistic about the role that artificial neural models will play in bridging the gap between theory and biology.
Of course, a number of outstanding issues remain: there is only so much we can learn from blackbox commercial models like ChatGPT that are trained on more data than any human could ever have access to. But we can train our own smaller models on carefully controlled data and then probe them using a wide variety of techniques. While it’s true many aspects of language models were developed for engineering ends, that doesn’t mean they can’t be repurposed for scientific research. Language models are overwhelmingly models of English and other major world languages, but that too is changing as models are trained on more and more languages. These are important issues, not insurmountable roadblocks.
@@
Baroni, (2022) lamented that, as of a 2021 exploration of citation records, Linzen et al.,’s seminal work on subject-verb agreeement was largely uncited within the linguistics community. But, since then, within the linguistics community, there are increasingly researchers using language models to ask questions within linguistics which are informed by and which can inform linguistic theory.
@@
Similarly, while there is a persistent stereotype that the NLP community cares mostly about engineering, we have found there to be significant interest in fundamentally linguistic questions. Of the 7 papers that won Best Paper Awards at ACL 2024, 1 was a direct response to a claim by Chomsky about language learning (Kallini et al.,, 2024), 1 was about inductive biases in models and of relevance to questions about constraints on language (Hahn and Rofin,, 2024), 1 was a theoretically motivated paper studying satisfiability in natural language and of relevance to old questions about the complexity of natural language (Madusanka et al.,, 2024), 1 was a new method for measuring memorization in models and is of relevance for studying trade-offs in memorization vs. generalization in natural language (Lesci et al.,, 2024), 2 were about reconstructing or recovering ancient languages (Lu et al.,, 2024; Guan et al.,, 2024), and 1 presented an open-access mulitilingual model for broadening coverage of under-resourced languages (Üstün et al.,, 2024). Every one of these papers has relevance to scientific questions about human language, and in most cases one or more of the authors works in a linguistics department and/or has a degree in linguistics.
@@
Thus, we don’t see ourselves as calling for radical change, but rather as acknowledging the interdisciplinary work that is already thriving in the 21st century. Linguistically informed computational work is increasingly taking place within linguistics departments, where computational researchers are working alongside syntacticians, semanticists, phonologists, language documentation experts, sociocultural linguistics, and experts in a wide variety of languages and language families. We think this is a very good development—one that both Norm and Claudette should embrace.
@@ 6 Acknowledgments
We thank Chris Potts for pointing out that he independently used a similar title in a talk called “Linguists for Deep Learning; or: How I Learned to Stop Worrying and Love Neural Networks” and encouraging us to use our title anyway. We thank Harvey Lederman for comments and helpful discussions about real patterns. We thank David Beaver, Ted Gibson, Greg Hickok, Laura Kalin, Connor Mayer, Kanishka Misra, Lisa Pearl, Greg Scontras, Steve Wechsler, and Xin Xie for helpful comments.
@@ 8 Competing Interest statement
The authors declare no competing interests.

### v2

@@ 4.7 Upshots for linguistics beyond language structure
In much of this piece, we have focused on the upshot of LMs for linguistic structure, particularly morphosyntax. We did so for two main reasons: first, because of the centrality of these topics within linguistics over the last 60+ years; and, second, because mastery of linguistic form emerged in models earlier than other high-level abilities (Mahowald et al.,, 2024). Taken together, these developments have led to important insights about human language learning and processing.
But LMs have also made recent striking gains in domains like reasoning, logic, and long-form dialog. These abilities seem to emerge not just from pretraining, but from techniques like supervised fine-tuning, instruction-tuning, and reinforcement learning from human feedback whereby models are given specific feedback to make them more aligned with human behavior on specific tasks (Ouyang et al.,, 2022; Bai et al.,, 2022; Achiam et al.,, 2023). The importance of inference-time compute—whereby models learn in-context and perform additional computation during the process of generating text—has also become apparent for tasks like math and reasoning (OpenAI et al.,, 2024; Marjanović et al.,, 2025).
As such, just as it is fruitful to explore what inductive biases and representational mechanisms underlie the emergence of syntax in LMs, it is increasingly fruitful to study what inductive biases and structures are necessary and/or sufficient for other abilities in LMs like reasoning (Marjanović et al.,, 2025), theory of mind (Hu et al.,, 2025), planning (Liu et al.,, 2023), and other aspects of higher-level cognition. Many of our same arguments hold in these domains as well: the upshot of the modern view of learning, the Contravariance Principle, the role of the statistical tradition, real patterns.
Moreover, as LMs continue to develop, they increasingly are becoming part of our speech communities. Millions of humans regularly interact with chatbots, in some cases carrying on long extended dialogs with chatbots playing various roles. And, increasingly, text read by humans is written or co-written by AI. How will the presence of AI change language? How should we think about dialogs between human and non-human interlocutors? What social consequences will emerge from this new paradigm? Linguistics as a field, particularly sociocultural linguistics and related disciplines, is well-positioned to be at the forefront of answering these questions (Bucholtz and Hall,, 2005; Meyerhoff,, 2006).
@@ 5 Conclusion
Rather than resisting this development, we contend it offers a golden opportunity for linguistics by enabling new kinds of research and opening up vistas of new hypotheses, methods, and research questions. Some of these questions will be new, like what kinds of artificial neural architectures best capture language, what kinds of biases they hold, and what the sources of those biases are. But we also have highlighted ways in which these methods can make progress on some of the oldest and most venerable questions in linguistics: what must be true of the input data for certain structures to be learned in principle, what are the constraints on what languages are possible, how does the form of language relate to its function.
In response to Piantadosi,’s claim that language models refute Chomsky, Kodner et al., (2023) wrote a piece skeptical of the role of LMs in linguistic theory called “Why linguistics will thrive in the 21st century”. We would go a step further and argue that, with the 21st century a quarter over, linguistics is thriving. But we think it’s doing so not in spite of language models, but in part because of them and the opportunities they open up for new research directions and the light they shed on old questions.
Therefore, we reject the dichotomy in the discourse around Piantadosi,’s (2023) paper, which often seems to pit language models against linguistics. Linguistic structure, as described in linguistic theories, is real and important even if language models learn those structures emergently in a complex statistical way. Conversely, language models can point the way to new ways of thinking about the learning, processing, and fundamental nature of language.
@@
  • Much of the field of linguistics has assumed that the form of language must be explained in terms of a symbolic formalism that constrains the forms of possible languages, and that this grammar formalism must represent innate human constraints on language, and that these constraints are logically necessary for language to be learned. The success of machine learning models should orient us away from this paradigm, because it shows that this approach is not necessary to characterize learning, and it opens up a new universe of statistical, quantitative, and functional theories to constrain the forms of possible languages and explain why only humans have language, while at the same time providing the tools to test those theories.
@@
  • Relatedly, the idea that the formal structure of linguistic competence should be the main focus of theoretical linguistics, which precedes other forms of study, is no longer tenable—language models reveal and allow us to characterize new dimensions of language such as how corpora of text express world knowledge, how the structure of usage supports learning or doesn’t, and how fundamental information-processing constraints shape the way that language is represented in brains and machines.
@@
  • The success of machine learning models shows us how language can be represented in ways that are graded, probabilistic, and fuzzy, which should move us away from an insistence on discrete categorical frameworks. They also reveal potential soft constraints on the structure of language.
@@
  • The ontological basis for linguistic structure need not lie in an innate genetic endowment. Linguistic structure is a real pattern, just as real and worthy of study whether it lies innate in the human genome or whether it is learned entirely through inductive statistical learning with domain-general biases. Linguistic phenomena do not become less interesting when they are learnable in this way, they become more interesting.
@@
More broadly, language models show that linguistics needs the rest of science. We will not answer the most fundamental questions about language by focusing only on language, using only methods that are tailor-made for linguistics. Just as neural networks revealed principles that are now fundamental in our understanding of vision, they are now in a position to revolutionize our understanding of language. Just as physics was revolutionized by thermodynamic ideas which arose from the ad-hoc analysis of complex machines, linguistics stands to benefit from new ways of thinking about computation arising from the analysis of neural networks. There is now massive intellectual activity spurred by language models in cognitive science, philosophy, physics, statistical learning theory, and information theory. The science of language can draw ideas, methods, and inspiration from all of this by maintaining a spirit of deep, curious, open-minded engagement and integration. Human language is in many ways unique as a natural phenomenon, and this means that the science of language should integrate with other fields of science more, in order to find deeper unities and to delineate and explain how and why language is special.
@@
In particular, we are optimistic about significantly revising what Stabler, (1983) said in doubting that “any close connections between linguistic theory and biology will be forthcoming”. He continues: “it is a crucial advantage of the computational approach that it has a functionalist vocabulary that does not require a type reduction to physicalistic concepts […] since this seems to be required in linguistics and in psychological accounts of language processing and visual perception”. We agree that there is a crucial advantage to the functionalist vocabulary used in linguistic theory. But by analogy to recent work in vision, we are optimistic about the role that artificial neural models will play in bridging the gap between theory and biology.
@@
This progress will come through an expansive linguistics — expansive in the breadth and diversity of languages it considers, expansive in its methods, and expansive in its connections to related fields. Baroni, (2022) lamented that, as of a 2021 exploration of citation records, Linzen et al.,’s seminal work on subject-verb agreement was largely uncited within the linguistics community. But, since then, there are increasingly researchers using language models to ask questions that are informed by and which can inform linguistic theory.
@@
On the computer science side, while there is a persistent stereotype that the Natural Language Processing (NLP) community cares mostly about engineering, we have found there to be significant interest in fundamentally linguistic questions. Of the 7 papers that won Best Paper Awards at ACL 2024 (the largest annual conference focused on NLP), 1 was a direct response to a claim by Chomsky about language learning (Kallini et al.,, 2024), 1 was about inductive biases in models of relevance to questions about constraints on language (Hahn and Rofin,, 2024), 1 was a theoretically motivated paper studying satisfiability in natural language of relevance to old questions about the complexity of language (Madusanka et al.,, 2024), 1 was a new method for measuring memorization in models and is of relevance for studying trade-offs in memorization vs. generalization in natural language (Lesci et al.,, 2024), 2 were about reconstructing or recovering ancient languages (Lu et al.,, 2024; Guan et al.,, 2024), and 1 presented an open-access mulitilingual model for broadening coverage of under-resourced languages (Üstün et al.,, 2024). Every one of these papers has relevance to scientific questions about human language, and in most cases one or more of the authors works in a linguistics department and/or has a degree in linguistics.
@@
Thus, we don’t see ourselves as calling for radical change, but rather as acknowledging the interdisciplinary work that is already thriving in the 21st century. Linguistically informed computational work is increasingly taking place within linguistics departments, where computational researchers are working alongside syntacticians, semanticists, phonologists, language documentation experts, sociocultural linguists, and experts in a wide variety of languages and language families. We think this is a very good development—one that both Norm and Claudette should embrace.
@@ 6 Acknowledgments
We thank Chris Potts for pointing out that he independently used a similar title in a talk called “Linguists for Deep Learning; or: How I Learned to Stop Worrying and Love Neural Networks” and encouraging us to use our title anyway. We thank David Beaver, Ted Gibson, Greg Hickok, Laura Kalin, Andy Kehler, Robbie Kubala, Harvey Lederman, Connor Mayer, Kanishka Misra, Andrea Moro, Lisa Pearl, Jon Rawski, Greg Scontras, Alex Warstadt, Steve Wechsler, Xin Xie, and three anonymous reviewers for helpful comments.
@@ 8 Competing Interest statement
KM has done contract work for OpenAI, unrelated to the content of this manuscript.